Three resistors of 10, 12 and "x" ohms, respectively are connected in parallel across a constant current source of 8A. Determine "x" if this resistor draws 2.5 A.


1
Expert's answer
2021-10-04T04:55:23-0400
"\frac{1}{10}+\frac{1}{12}+\frac{1}{x}=\frac{1}{2.5\cdot x}8\\x=12\Omega"